# Quantum Delocalization Enables Water Dissociation on Ru(0001)

### v2

@@ Abstract
Having identified the most stable intact and dissociative overlayers, we now turn to tackling the enduring question of whether water overlayers dissociate on Ru(0001). First, we computed the dissociation energy barrier, yielding 0.53 eV from the H-down to the half-dissociated model, aligning with previous predictions Michaelides et al. (2003a); Meng et al. (2005). The energy barrier is found to be higher (0.65 eV) for the extended chains model to dissociate an H2O (Supplemental Material Fig. S5 SM ). Based on a rough estimation from transition state theory Truhlar et al. (1996), a dissociation energy barrier of this magnitude would necessitate a time scale exceeding microseconds, far beyond the accessible time-scale of MD simulations. Indeed, we did not observe the dissociation throughout a 0.5 $\mu$s long classical MD simulation [Fig. 2(a)]. Nevertheless, classical MD simulations do not take into account the NQEs, which have been shown to be significant in hydrogen-related systems Marx and Parrinello (1996); Benoit et al. (1998); Marx et al. (1999); Kumagai et al. (2008); Morrone et al. (2009); Li et al. (2010, 2011); Chen et al. (2013); Feng et al. (2015); Guo et al. (2016); Ceriotti et al. (2016); Guo et al. (2017); Chen et al. (2018); Markland and Ceriotti (2018); Wei Fang and Michaelides (2019); Cheng et al. (2019); Lan et al. (2020); Bore and Paesani (2023); Flór et al. (2024). In particular, Li et al. Li et al. (2010) showcased the pronounced NQEs for water overlayers on transition metal surfaces through ab initio PIMD simulations. Despite their high accuracy, ab initio methods face limitations concerning the accessible time and length scales, hindering the direct observation of water dissociations.
@@
Thanks to our accurate and efficient machine-learning potential, we were able to perform PIMD simulations on large supercells for an extended timescale, while maintaining ab initio accuracy. In the PIMD simulations, 16 beads were employed to sample the imaginary time path-integral and the statistical analyses were derived by averaging these 16 beads. We employed a $6\times 4\sqrt{3}$ supercell with 32 H2O molecules based on the most stable intact overlayer configuration (i.e., extended chains model) as the initial structure. Throughout a 60 ns PIMD simulation, two abrupt changes in the number of OH, H2O, and H3O species are observed [see Fig. 2(b)], indicating two instances of water dissociation. To corroborate our observation, we conducted additional independent PIMD and classical MD simulations. The water dissociation is reproduced in every PIMD simulation, but is not observed in any of the classical MD simulations within the considered timescale (Supplemental Material Fig. S6 SM ). These results demonstrate that the NQEs facilitate the water dissociation on Ru(0001).
@@
To unveil the details of proton transfers and water dissociation, we analyzed the PIMD trajectory. Note that although PIMD does not provide a real-time picture of the trajectories, we still believe that the mechanisms we observe would prevail in a centroid molecular dynamics and that the observed mechanisms are highly relevant to water dissociation. Initially, ephemeral hydroxyls exist due to the quantum delocalization associated with the NQEs, as evidenced by the fluctuation between the number of hydroxyls and intact H2O (Supplemental Material Fig. S8 SM ). At 3.8975 ns, the formation of an H3O and OH pair commences due to the proton transfer [Figs. 2(f)-(g)], and the pair would typically last for $\sim$38 ps (Supplemental Material Fig. S8 SM ). At 3.9355 ns, the first instance of water dissociation emerges. Interestingly, we observe that the dissociation does not directly originate from H3O, but instead undergoes an indirect process. Specifically, one of the protons in H3O is transferred to the neighboring H2O, leading to the simultaneous dissociation of a hydrogen atom from the latter [Figs. 2(h)-(i)]. The dissociated H is observed to preferentially occupy the hollow fcc sites rather than the top sites on Ru(0001). This observation is consistent with our DFT calculations, when the dissociation fraction is low [Supplemental Material Fig. S5(f) SM ]. Moreover, we find that once dissociated, the dissociated H decreases its vertical height [Fig. 2(e)], diffuses on the surface, but is rarely involved in proton transfers. The existence of the dissociated hydroxyl increases the frequency of proton transfers [Fig. 2(c)]. At 29.1005 ns, the second occurrence of water dissociation takes place, following a procedure akin to the first instance [Figs. 2(j)-(k) and Supplemental Material Fig. S9 SM ]. At this point, the system contains 30 intact water molecules, 2 hydroxyls, and 2 dissociated hydrogen atoms. Over an extended time scale, a larger number of H2O would likely undergo dissociation. However, the large computational cost of the PIMD kept us from much longer simulations.
@@
In summary, we have developed an efficient and robust machine-learning potential to accurately capture the intricate interactions between water overlayers and the Ru(0001) surface through active learning and basis sets optimization. This enables us to identify the most stable overlayers for both intact and dissociative adsorption and facilitate the execution of PIMD simulations with near first-principles accuracy on a time and length scale that is inaccessible through ab initio methods. This capability is crucial, as it allows us to directly observe water dissociation on Ru(0001) through a nanosecond-scale PIMD simulation, a remarkable achievement that is unattainable with previous ab initio studies. We clearly demonstrate that the water dissociation is driven by the quantum delocalization originating from NQEs. These are essential to facilitate rapid and frequent proton transfers. Furthermore, the formation of H3O and OH pairs appears to occur via long-range density fluctuations, with an H3O and OH pair forming quasi-instantly over a distance of up to 5 connected H2O molecules. This is facilitated by a combination of the Grotthuss mechanism and quantum fluctuations. To our knowledge, such a long-range mechanism has not been observed before. Our study implies that quantum fluctuations can dramatically change the barriers to water dissociation, but also lead to new dissociation pathways. This has far-reaching implications not only for surface science, electrolysis and fuel cells, but also for biology, where water is ubiquitous. Clearly, our study suggests that quantum fluctuations must be taken into account whenever water dissociation occurs. The implications for the now so important field of hydrogen production are huge, as the barriers appear to be cut in half by quantum fluctuations.
@@
This work is supported by the National Natural Science Foundation of China (Grants No. 52422112, No. 52188101, and No. 52201030), the Liaoning Province Science and Technology Planning Project (2023021207-JH26/103 and RC230958), the National Key R&D Program of China 2021YFB3501503, and the Special Projects of the Central Government in Guidance of Local Science and Technology Development (2024010859-JH6/1006).

### v3

@@ Abstract
Having identified the most stable intact and dissociative overlayers, we now turn to tackling the enduring question of whether water overlayers dissociate on Ru(0001). First, we computed the dissociation energy barrier, yielding 0.53 eV for dissociating half of the water molecules, aligning with previous predictions Michaelides et al. (2003a); Meng et al. (2005). The energy barrier is found to be higher ($>$0.6 eV) for dissociating 1/8 of the water molecules (Supplemental Material Fig. S5 SM ). Based on a rough estimation from transition state theory Truhlar et al. (1996), a dissociation energy barrier of this magnitude would necessitate a time scale exceeding microseconds, far beyond the accessible time-scale of MD simulations. Indeed, we did not observe the dissociation throughout a one-microsecond classical MD simulation [Fig. 2(a)]. Nevertheless, classical MD simulations do not take into account the NQEs, which have been shown to be significant in hydrogen-related systems Marx and Parrinello (1996); Benoit et al. (1998); Marx et al. (1999); Kumagai et al. (2008); Morrone et al. (2009); Li et al. (2010, 2011); Chen et al. (2013); Feng et al. (2015); Guo et al. (2016); Ceriotti et al. (2016); Guo et al. (2017); Chen et al. (2018); Markland and Ceriotti (2018); Fang et al. (2019); Cheng et al. (2019); Lan et al. (2020); Bore and Paesani (2023); Flór et al. (2024). In particular, Li et al. Li et al. (2010) showcased the pronounced NQEs for water overlayers on transition metal surfaces through ab initio PIMD simulations. Despite their high accuracy, ab initio methods face limitations concerning the accessible time and length scales, hindering the direct observation of water dissociations.
@@
Thanks to our accurate and efficient machine-learning potential, we were able to perform PIMD simulations on large supercells for an extended timescale, while maintaining ab initio accuracy. In the PIMD simulations, 16 beads were employed to sample the imaginary time path-integral and the statistical analyses were derived by averaging these 16 beads. We employed a $6\times 4\sqrt{3}$ supercell with 32 H2O molecules based on the most stable intact overlayer configuration (i.e., extended chains model) as the initial structure. Throughout a 200-nanosecond PIMD simulation, three abrupt changes in the number of OH, H2O, and H3O species are observed [see Fig. 2(b)], indicating three instances of water dissociation. To corroborate our findings, we conducted additional independent PIMD and classical MD simulations. Water dissociation was consistently reproduced in all PIMD simulations but was absent in classical MD simulations within the considered timescale (Supplemental Material Fig. S7 SM ). Furthermore, PIMD simulations of D2O adsorption on Ru(0001), conducted under identical computational settings as H2O adsorption, revealed no dissociation. This result underscores the pronounced kinetic isotope effect and aligns with experimental observations Clay et al. (2004). Collectively, these results demonstrate that the NQEs facilitate the water dissociation on Ru(0001).
@@
To unveil the details of proton transfers and water dissociation, we analyzed the PIMD trajectory. Note that although PIMD does not provide a real-time picture of the trajectories, we still believe that the mechanisms we observe would prevail in a centroid molecular dynamics and that the observed mechanisms are highly relevant to water dissociation. Initially, ephemeral hydroxyls exist due to the quantum delocalization associated with the NQEs, as evidenced by the fluctuation between the number of hydroxyls and intact H2O (Supplemental Material Fig. S9 SM ). At 3.8975 ns, the formation of an H3O and OH pair commences due to the proton transfer [Figs. 2(f)-(g)], and the pair would typically last for $\sim$38 ps (Supplemental Material Fig. S9 SM ). At 3.9355 ns, the first instance of water dissociation emerges. Interestingly, we observe that the dissociation does not directly originate from H3O, but instead undergoes an indirect process. Specifically, one of the protons in H3O is transferred to the neighboring H2O, leading to the simultaneous dissociation of a hydrogen atom from the latter [Figs. 2(h)-(i)]. The dissociated H is observed to preferentially occupy the hollow fcc sites rather than the top sites on Ru(0001). This observation is consistent with our DFT calculations, when the dissociation fraction is low [Supplemental Material Fig. S5(l) SM ]. Moreover, we find that once dissociated, the dissociated H decreases its vertical height [Fig. 2(e)], diffuses on the surface, but is rarely involved in proton transfers. The existence of the dissociated hydroxyl increases the frequency of proton transfers [Fig. 2(c)]. At 29.1005 ns, the second occurrence of water dissociation takes place, following a procedure akin to the first instance [Figs. 2(j)-(k) and Supplemental Material Fig. S10 SM ]. At 179.9685 ns, a third water molecule dissociates [Fig. 2(b)]. Notably, this third dissociation event does not involve the formation of H3O (Supplemental Material Fig. S11 SM ). At this point, the system contains 29 intact water molecules, 3 hydroxyls, and 3 dissociated hydrogen atoms. Over extended time scales, a greater number of H2O would likely dissociate. Note that within the simulated time frame, no water desorption was observed due to the higher energy barrier for desorption compared to dissociation (Supplemental Material Figs. S6 and S5 SM ). This aligns with experimental findings that H2O dissociates prior to desorption Clay et al. (2004). However, the large computational cost of PIMD kept us from much longer simulations.
@@
In summary, we have developed an efficient and robust machine-learning potential to accurately capture the intricate interactions between water overlayers and the Ru(0001) surface through active learning and basis sets optimization. This enables us to identify the most stable overlayers for both intact and dissociative adsorption and facilitate the execution of PIMD simulations with near first-principles accuracy on a time and length scale that is inaccessible through ab initio methods. This capability is crucial, as it allows us to directly observe water dissociation on Ru(0001) through a nanosecond-scale PIMD simulation, a remarkable achievement that is unattainable with previous ab initio studies. We clearly demonstrate that the water dissociation is driven by the quantum delocalization originating from NQEs. These are essential to facilitate rapid and frequent proton transfers. Furthermore, the formation of H and OH pairs appears to occur via long-range density fluctuations, with an H and OH pair forming quasi-instantly over a distance of up to 5 connected H2O molecules. This is facilitated by a combination of the Grotthuss mechanism and quantum fluctuations. To our knowledge, such a long-range mechanism has not been observed before. Our study implies that quantum fluctuations can dramatically change the barriers to water dissociation, but also lead to new dissociation pathways. This has far-reaching implications not only for surface science, electrolysis and fuel cells, but also for biology, where water is ubiquitous. Clearly, our study suggests that quantum fluctuations must be taken into account whenever water dissociation occurs. The implications for the now so important field of hydrogen production are huge, as the barriers appear to be cut in half by quantum fluctuations.
@@
This work is supported by the National Natural Science Foundation of China (Grants No. 52422112, No. 52188101, and No. 52201030), the Strategic Priority Research Program of the Chinese Academy of Sciences (XDA041040402), the Liaoning Province Science and Technology Major Project (2024JH1/11700032, 2023021207-JH26/103 and RC230958), the National Key R&D Program of China 2021YFB3501503, and the Special Projects of the Central Government in Guidance of Local Science and Technology Development (2024010859-JH6/1006).
Data availability—The data that support the findings of this article are openly available Dat .